OpenStudy (clarence):

Only started the topic today and I don't quite understand it. :/ For example, how do you determine the stationary points and point of inflection for y = x^3 - 3x^2 - 9x + 7

OpenStudy (anonymous):

Stationary points are the points the rate of change of the function is zero. It means the places the value of the function changes the least as you vary x. Try to calculate the value of \(\Delta y=f(x+\Delta x)-f(x)\). It will help you to understand it. Also you can go to the concept of the derivative and understand it intuitively. There is a great lesson on that over at the khanacademy you should watch it: https://www.khanacademy.org/math/calculus/differential-calculus/derivative_intro/v/calculus--derivatives-1--new-hd-version

OpenStudy (anonymous):

\[\Delta y=f(x+\Delta x)-f(x)\\ \Delta y=(x+\Delta x)^3-3(x+\Delta x)^2-9(x+\Delta x) +7 -(x^3-3x^2-9x +7) \] If you go ahead and try to solve this, you will notice that \(\Delta y\) is a function of both \(\Delta x\) and x, and there should be places where it changes very slowly.

OpenStudy (clarence):

Wait, so how would I have solved y = x^3-3x^2-9x+7 with that?

OpenStudy (anonymous):

No, what I meant was to think about it. What are the values of \(x\) which \(\Delta y\) changes the least? That is a stationary point

OpenStudy (clarence):

Oh, okay, I kinda get it. :p

OpenStudy (anonymous):

Can you graph that equation? Then I can show to you on the graph. Its much simpler.

OpenStudy (clarence):

How do I graph it on here?

terenzreignz (terenzreignz):

Can you differentiate?

OpenStudy (clarence):

Yep! :)

terenzreignz (terenzreignz):

Okay, so can you differentiate the function?

terenzreignz (terenzreignz):

\[\large y = x^3 - 3x^2 - 9x+8\] differentiate... \[\Large \frac{dy}{dx} = ?\]

OpenStudy (dan815):

hello by stationary if u mean critial points then it is asking you to find the the points where first derivative =0 and for points of inflection where 2nd derivative = 0

OpenStudy (dan815):

the first derivative gives you a slope graph, so it is telling you the rate of change at every point for there to be a "stationary" point the rate of change must = 0 so the 1st derivative must = 0

OpenStudy (dan815):

an inflection point happens when the graphs concavity is changing or you can think of this as acceleration to deceleration and viceversa

OpenStudy (dan815):

or the rate of change for the rate of change is 0, because if the rate of change of the rate of change is going from positive to negative or negative to positive it must go thru the value of 0, so we calculate the 0 point of the 2nd derivative, and see how its changing to either side of the zero to determine if its a max or a min, kind of a lot to take in i know xD

OpenStudy (clarence):

Aha, yeah, Math's Studies is harder than I thought it would be. :p

OpenStudy (dan815):

just keep at it, this is where math gets fun :)

OpenStudy (dan815):

take one step at a time, first do you understand what the first derivative is

terenzreignz (terenzreignz):

Differentiate y... have you done it yet?

OpenStudy (dan815):

oh i just realized the first guy posted you a khan academy link... did u watch that

OpenStudy (clarence):

Internet's slow so I can't really watch stuff :/ But yes I do know what the first derivative is :) and differentiating y I'm pretty sure you'll get y = 3x^2 - 6x -9? :)

terenzreignz (terenzreignz):

okay, work with that... now you can equate that to zero.. \[\large 3x^2 - 6x - 9 = 0\ and solve for x

terenzreignz (terenzreignz):

\[\large 3x^2 - 6x - 9 = 0\]